### v1

@@ 6 Conclusion
In this study, we propose an effective-Rank based Feature Richness (RFR) enhancement method that can be integrated with a wide range of existing methods in CIL. More specifically, our method increases the effective rank of representations during the base session. In order to substantiate the effectiveness of our method, we have established a theoretical connection between the effective rank and the Shannon entropy of representations. Through empirical analysis, we have demonstrated the efficacy of our method in several dimensions including enhancement in forward compatibility, mitigation of catastrophic forgetting, and improvement in performance. In summary, our method effectively enhances the performance by increasing effective rank and feature richness in the representation.

### v2

@@ 6 Conclusion
In this study, we propose effective-Rank based Feature Richness (RFR) enhancement method that can be integrated with a wide range of existing methods in CIL. More specifically, our method increases the effective rank of representations during the base session. In order to substantiate the effectiveness of our method, we have established a theoretical connection between the effective rank and the Shannon entropy of representations. Through empirical analysis, we have demonstrated the efficacy of our method in several dimensions including enhancement in forward compatibility, mitigation of catastrophic forgetting, and improvement in performance. In summary, our method effectively enhances both forward and backward compatibility simultaneously. The superiority of this method is supported by both theoretical and empirical evidences.